OpenStudy (anonymous):

Simplify the radical. √36a^4b^10

OpenStudy (anonymous):

\[\sqrt{36*a^{4b ^{10}}}=\] is your expression?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its

OpenStudy (austinl):

\(\sqrt{36\times a^4\times b^{10}}\) I assume.

OpenStudy (anonymous):

sqrt36=6 sqrta^4=a^2 sqrtb^10=b^5

OpenStudy (anonymous):

multiply them all

OpenStudy (austinl):

That is, \(\sqrt{36}\times \sqrt{a^4}\times\sqrt{b^{10}}=6\times a^2\times b^5=6a^2b^5\)

OpenStudy (anonymous):

okay, so this is √a^3

OpenStudy (anonymous):

a?